# NoTeeline: Supporting Real-Time, Personalized Notetaking with LLM-Enhanced Micronotes

### v2

@@ Appendix A Participants’ Information
ID Gender Occupation
P1 Male Graduate Student
P2 Female Undergraduate
P3 Female UX Designer
P4 Female Undergraduate
P5 Female Masters Student
P6 Female Student
P7 Female Graduate student
P8 Female Masters Student
P9 Male Undergraduate
P10 Female Student
P11 Male Product Manager
P12 Male Software Engineer
Figure 12. Demography of user study participants.
@@ Appendix C Relevance of expanded note and summary
Appendix D Prompts used for NoTeeline
D.1. Prompt for Note Expansion
@@
D.2. Prompt for Note Organization by Theme
@@ D.4. Prompt for Summary
Appendix E Questionnaire for NeedFinding Study
@@
{mdframed}
[] I want you to act as a personalized note-taking assistant. Users will give you a keypoint and the YouTube transcript. Your task is to expand the keypoint into a note point, by taking additional context from the transcript. The note should be a full sentence in simple English. Follow these rules:
@@
{mdframed}
[] Given a list of points, Your task is to perform topic modeling over them. Arrange the points into topics and provide a name for each topic. Please mark the topic within <Topic></Topic>tags. the points are marked with <p>tag as well. Follow these rules:
@@
{mdframed}
[] System prompt:
@@
{mdframed}
[] I will give you a context and some keypoints, Your task is to summarize the keypoints in 4 sentences. Focus on the keypoint, only use context if you need extra information:
Context: Preprocessed Transcript in Langchain
Keypoints: Expanded Points
Figure 16. Prompt for Summary
P1 P2 P3 P4 P5 P6 P7 P8 P9 P10 P11 P12

### v3

@@ Appendix A Participants’ Information
[…]
Different tools and techniques participants reported to use for note-taking in their daily lives.
@@ Appendix C Relevance of expanded note and summary
Appendix D Response Time for OpenAI API
Table 8 shows the response time for the OpenAI API. On average, OpenAI takes $2$ seconds to generate the expanded notes from micronotes.
ID
API Response Time
(in seconds)
P1 1.83
P2 1.47
P3 1.53
P4 1.43
P5 2.25
P6 1.75
P7 1.71
P8 2.67
P9 2.61
P10 2.09
P11 2.37
P12 2.38
avg. 2.00
Table 8. API response time (in seconds) for each participant.
Appendix E Prompts used for NoTeeline
E.1. Prompt for Note Expansion
@@
E.2. Prompt for Note Organization by Theme
@@ E.4. Prompt for Summary
Appendix F Questionnaire for Needfinding Study
@@
I want you to act as a personalized note-taking assistant. Users will give you a keypoint and the YouTube transcript. Your task is to expand the keypoint into a note point, by taking additional context from the transcript. The note should be a full sentence in simple English. Follow these rules:
@@
[…]
Prompt used for note expansion
Given a list of points, Your task is to perform topic modeling over them. Arrange the points into topics and provide a name for each topic. Please mark the topic within <Topic></Topic>tags. the points are marked with <p>tag as well. Follow these rules:
@@
[…]
Prompt for Note Organization by Theme
I will give you a context and some keypoints, Your task is to summarize the keypoints in 4 sentences. Focus on the keypoint, only use context if you need extra information:
Context: Preprocessed Transcript in Langchain
Keypoints: Expanded Points
Figure 14. Prompt for Summary
[…]
Prompt for Summary
System prompt:
@@
[…]
Prompt for Cue Question